### v1

# RF-MatID: Dataset and Benchmark for Radio Frequency Material Identification

Abstract
Accurate material identification plays a crucial role in embodied AI systems, enabling a wide range of applications. However, current vision-based solutions are limited by the inherent constraints of optical sensors, while radio-frequency (RF) approaches, which can reveal intrinsic material properties, have received growing attention. Despite this progress, RF-based material identification remains hindered by the lack of large-scale public datasets and the limited benchmarking of learning-based approaches. In this work, we present RF-MatID, the first open-source, large-scale, wide-band, and geometry-diverse RF dataset for fine-grained material identification. RF-MatID includes 16 fine-grained categories grouped into 5 superclasses, spanning a broad frequency range from 4 to 43.5 GHz, and comprises 142k samples in both frequency- and time-domain representations. The dataset systematically incorporates controlled geometry perturbations, including variations in incidence angle and stand-off distance. We further establish a multi-setting, multi-protocol benchmark by evaluating state-of-the-art deep learning models, assessing both in-distribution performance and out-of-distribution robustness under cross-angle and cross-distance shifts. The 5 frequency-allocation protocols enable systematic frequency- and region-level analysis, thereby facilitating real-world deployment. RF-MatID aims to enable reproducible research, accelerate algorithmic advancement, foster cross-domain robustness, and support the development of real-world application in RF-based material identification.
[…]
1 Introduction
Material identification is the task of identifying an object’s physical category (e.g., metal, plastic) from its intrinsic properties, which is fundamental for autonomous systems, particularly in embodied intelligence. Embodied intelligence arises from an agent’s physical interaction with the environment, enabling it to perceive, reason, and adapt. Accurate material identification allows an embodied agent to infer object properties within its operational range, guiding manipulation and functional interactions. This capability has broad implications across diverse domains, such as enabling embodied agents to interpret fine-grained object attributes for physical scene understanding through material perception (xiao2018unified), and to ground complex reasoning and control in material-driven functional affordances (do2018affordancenet)
Currently, material identification relies mainly on optical sensors, such as cameras and hyperspectral sensors (drehwald2023one; xue2017differential; schmid2023material), to capture distinguishable spatial characteristics (e.g. texture, edges) (erickson2020multimodal) and light spectrum features (e.g. reflectance and transmittance values) (zahiri2022comparison). However, vision-based material identification is inherently constrained by the visual similarity of fine-grained categories (e.g., steel vs. aluminum), limited robustness under real-world perturbations such as lighting and perspectives, and the inability of sensors to reveal intrinsic physical properties, including elasticity and conductivity.
To overcome the inherent constraints of optical sensors, non-visual sensing modalities, such as radio frequency (RF), are gaining attention as they exploit electromagnetic interactions to reveal intrinsic material properties beyond surface appearance (khushaba2022radar; hagele2025occluded). However, RF-based material identification has not yet been extensively explored. There are no publicly accessible large-scale datasets, hindering reproducibility and fair benchmarking across algorithms. In addition, commercial off-the-shelf (COTS) sensors offer limited and fragmented frequency coverage, hindering systematic evaluation of material identification across the RF spectrum, which is crucial for selecting optimal operational bands in various applications. Finally, most studies are conducted in controlled laboratory settings and rarely incorporate perturbations such as variations in sensor–object geometry (e.g., incidence angle, stand-off-distance), leaving open questions regarding robustness and scalability in deployment scenarios.
Thus, we present RF-MatID, a large-scale, wide-band, and geometry-diverse RF dataset designed to advance fine-grained material identification and enable the development of more robust algorithms. RF-MatID encompasses 16 carefully curated fine-grained categories, systematically derived from 5 superclasses to capture subtle intra-class variations. The dataset spans a broad frequency spectrum from 4 to 43.5 GHz, sampled at 53 points per GHz. Through large-scale data acquisition, RF-MatID comprises 142k samples represented in both the frequency and time domains (71k samples in each). Note that these 142k samples correspond to representation-level instances derived from 71k unique physical measurements, each provided with a paired frequency- and time-domain representation. RF-MatID also provides 5 frequency protocols (including protocols compliant with legal frequency regulations in major global economies) and 7 split settings for a comprehensive benchmark. The key contributions and characteristics of RF-MatID are summarized as follows:
  • First open-source, wide-band, geometry-diverse RF dataset for fine-grained material identification. To the best of our knowledge, we construct the first large-scale open-source RF dataset covering 16 fine-grained material categories from 5 superclasses. The dataset spans a wide frequency band (4–43.5 GHz), and systematically incorporates variations in incidence angle and stand-off distance to emulate realistic geometric conditions. Both time-domain and frequency-domain representations are provided, enabling multi-perspective evaluation.
  • Investigation of RF data representations and protocol-level applicability. We evaluate RF data representations by comparing classification accuracy on raw frequency- and processed time-domain signals, showing that raw frequency-domain data can be directly leveraged by deep learning models without additional domain transformation. Additionally, we explore RF-based material identification under various frequency-allocation protocols to assess the practicality of deploying RF systems in compliance with regulatory constraints. We further evaluate consecutive sub-bands of different bandwidths to gain insights into frequency selection across diverse applications.
  • Benchmark of learning-based approaches and robustness. We establish extensive benchmarks of state-of-the-art deep learning models on the RF material identification task, adapting architectures from computer vision and natural language processing. Beyond in-distribution accuracy, we present a systematic evaluation of out-of-distribution robustness in RF sensing, through cross-angle and cross-distance domain shifts that emulate geometric perturbations.
2 Related works
2.1 Material Identification
Material identification has been extensively studied due to its importance in industrial (johns2023framework) and civil (ha2018learning) contexts in the previous decades. In the signal processing domain, physics-based approaches have been well studied for material identification (wolff1990polarization; wu2020msense; rothwell2016analysis; sahin2020simplified), which employ signal processing techniques such as the Fresnel equations (fresnel1834memoire) and the Nicolson–Ross–Weir (NRW) method (nicolson2007measurement) to estimate intrinsic material electromagnetic properties (e.g. permittivity, conductivity, and absorption characteristics) for materials distinguishing. However, physics-based approaches often rely on idealized assumptions and hand-crafted formulas, introducing challenges such as sensitivity to environmental conditions, limited adaptability to diverse application scenarios, and susceptibility to extraneous object-specific factors unrelated to the material(e.g. thickness and size).
Dataset Modality Frequency (GHz) Band- Width # Super- Classes # Sub- Classes Geometric Variations # Samples Public Accessibility # Benchmark Models
ha2020food RFID 0.5-1.0 0.5 1 16 distance, angle 2,048 4
wang2017tagscan 0.92-0.93 0.01 2 16 distance, angle 1,800 -
feng2019wimi Wi-Fi 5.0-5.8 0.8 1 10 distance 2,000 -
shi2021environment 2.4-2.5 0.1 4 14 - 1,568 -
dhekne2018liquid UWB 3.5-5.5 2.0 1 33 - 330 -
zheng2021siwa 6.5-8.0 1.5 2 4 distance 14,000 3
wu2020msense mmWave 58.7-61.3 2.6 5 21 distance 67,200 -
he2022accurate 77.0-81.0 4.0 5 - angle 200,000 2
shanbhag2023contactless 77.0-81.0 4.0 7 23 distance, angle 15,386 3
chen2025material 77.0-81.0 4.0 6 - distance, angle 129,600 5
RF-MatID UWB-mmWave 4.0-43.5 39.5 5 16 distance, angle 142,000 9
Table 1: Comparisons of RF-MatID with other RF material identification datasets. Previous datasets are organized by sensing modality and a ‘’ in the table denotes the absence of corresponding dataset feature in the research work.
With the advancement of deep learning techniques, learning-based approaches provide more accurate and generalizable solutions for material identification. Vision-based approaches constitute the predominant paradigm for material identification. These methods exploit visual features, such as color, texture, and reflectance properties, allowing computer vision models to learn hierarchical representations that reveal the underlying material patterns. Driven by the ubiquitous deployment of optical sensors, numerous material image datasets and benchmarks have been established (weinmann2014material; bell2015material), enabling standardized and reproducible research and improving model accuracy and generalizability (drehwald2023one; xue2017differential). Beyond RGB-based solutions, other studies have explored more physically informative optical modalities: for instance, TOF cameras capture a combination of surface and subsurface scattering effects (su2016material), while hyperspectral imaging records a full spectral profile at each image pixel (salas2025hyperspectral). However, the weak penetration of near-infrared camera signals, which are typically in the MHz range, limits their ability to capture information beyond an object’s surface. As a result, vision-based material identification is highly susceptible to variations in lighting and object geometry, which has spurred greater interest in robust RF-based approaches.
Existing RF-based deep learning approaches for material identification can be broadly divided into feature-based (two-stage) methods and end-to-end learning methods. Feature-based methods extend traditional physics-driven approaches by applying machine learning classifiers on manually engineered signal features for material identification (he2022accurate; shanbhag2023contactless). While such classifiers can capture latent feature patterns beyond hand-crafted rules, their performance remains constrained by the quality of engineered features and shows limited adaptability to diverse real-world conditions. To overcome the limitations, recent works explore end-to-end deep learning frameworks that directly operate on RF signals, enabling discriminative representation learning of material characteristics without explicit feature extraction. (zhang2024airtac; hagele2025occluded).
2.2 Radio-Frequency Material Datasets
Currently, most RF-based material datasets are collected using wireless COTS sensors operating in the 0.9–81 GHz frequency range, capturing electromagnetic signals that encode material properties (chen2024wireless). Accordingly, these datasets can be broadly categorized based on the sensing modality into four clusters. Radio-Frequency-Identification (RFID)-based datasets are mostly collected in the 0.90–0.927 GHz frequency band, and the target object is either tagged on the surface or placed between a reader and tags deployed in the environment (wang2017tagscan; ha2020food). The modality enables applications such as through-wall sensing via strong low-frequency penetration, but the reliance on physical tags greatly complicates data collection. Wi-Fi-based datasets primarily leverage the 2.4 GHz and 5 GHz frequency bands, and the target object is typically placed between a Wi-Fi transmitter and receiver (feng2019wimi; shi2021environment). The ubiquity of Wi-Fi devices facilitates convenient data collection, but the narrow bandwidth and distortions induced by hardware imperfections constrain the reliability of signals. Ultra-WideBand (UWB)-based datasets are mostly collected using ultra-wideband pulse signals in the 3.1–10.6 GHz frequency range and the equipment setup is similar to WiFi-based collection. UWB-based systems can capture fine-grained attenuation and delay properties (dhekne2018liquid; zheng2021siwa), but face practical barriers from strict synchronization demands and high deployment costs. Millimeter-Wave (MmWave)-based datasets are mainly collected using COTS devices operating in the 57–64 GHz and 76–81 GHz frequency bands (wu2020msense; shanbhag2023contactless; chen2025material). A single mmWave radar unit is an integrated transceiver that provides accurate phase information characterizing the dielectric properties of the materials. However, its limited penetration confines the sensing to material surfaces, making it highly sensitive to environmental occlusions.
Figure 1: Data collection setup: (a) the customized sensing platform and (b) the acquisition layout.
However, we observe that none of the existing datasets are publicly accessible, resulting in a lack of standardized state-of-the-art comparisons across the latest algorithms. Moreover, due to limitations of COTS sensors, each dataset typically covers only a narrow frequency band, lacking diversity in the frequency domain. In addition, most datasets lack systematic benchmarks to evaluate the performance of modern deep learning models on RF-based material identification tasks. To address these gaps, we propose RF-MatID, the first open-source dataset and benchmark covering 39.5 GHz UWB-mmWave (zhang2013background) frequency band, including 16 fine-grained material categories, and providing comprehensive benchmarks with 9 models, 5 evaluation protocols, and 7 data split settings. RF-MatID is critical for advancing machine learning research in RF-based material identification and can facilitate developments in embodied AI for tasks such as indoor scene understanding, precise robotic manipulation, and affordance learning. Table 1 summarizes both previously used RF-based material datasets and our proposed RF-MatID dataset.
3 Preliminaries of RF Sensing
3.1 RF data properties
Various RF-based material sensing systems, summarized in section 2.2, can be broadly categorized into radar-based and non-radar-based approaches. Benefiting from coherent transceiver design, radar-based sensing provides high-resolution amplitude and phase information, that can serve as discriminative features for learning-based material classification in indoor embodied AI tasks. Leveraging these advantages, we establish a UWB-mmWave sensing platform designed to drive real-world applications. In RF-MatID’s mono-static sensing system, electromagnetic waves are transmitted from an antenna, interact with the material subject, and reflect back to the antenna. When these waves encounter a material, a portion of their energy is reflected at the surface, while the remainder is transmitted into the material. The material’s intrinsic physical properties, such as permittivity and conductivity, will affect the amplitude, phase, and temporal characteristics of the electromagnetic waves. These variations form informative latent features for machine learning models, enabling fine-grained material identification. Furthermore, the behavior of electromagnetic waves is influenced by its frequency: lower-frequency waves penetrate deeper and reveal bulk properties, whereas higher-frequency waves are more sensitive to surface details but attenuate more in lossy materials. Thus, our system collects signal data spanning both the centimeter-wave band (3–30 GHz) and the millimeter-wave Q-band (30–50 GHz), ensuring complementary information capture for robust learning across diverse materials.
In our RF-based material sensing setup, the acquired data consists of complex signals uniformly sampled across the 4–43.5 GHz band. The spectrum is discretized into 2,048 frequency bins, each corresponding to a distinct carrier frequency. As illustrated in the formula below, the response at each frequency $f_{i}$ is represented by its in-phase ($I$) and quadrature ($Q$) components, forming a complex value.
$H(f_{i})=I(f_{i})+jQ(f_{i})=\lvert H(f_{i})\rvert\,e^{j\angle H(f_{i})}$ (1)
Here, the magnitude $\lvert H(f_{i})\rvert$ encodes amplitude attenuation and the phase $\angle H(f_{i})$ encodes the propagation delay introduced by the material. These complementary features form the raw sensor measurements, but standard deep learning backbones operate in the real domain. This discrepancy motivated the exploration of efficient RF representations, as presented in section 4.2, enabling learning-based models to effectively leverage complementary amplitude and phase information for fine-grained material identification.
3.2 RF tools and platform
To facilitate the data acquisition of our RF-MatID, we develop a customized RF data collection platform. As shown in Figure 1(a), we employ an RF SPIN DRH40 (drh40) double ridged horn antenna to transmit and receive signals across 4–40 GHz. The signals are subsequently processed by a 1-port vector network analyzer (MS46131A) (ms46131a), operating over 1 MHz–43.5 GHz, to generate the raw frequency-domain data. The sensor’s sensing range is $\sim$2 m, which is intentionally tailored for indoor robot manipulation tasks in embodied AI that require high-precision, close-range perception. This design choice facilitates applications where the sensor is mounted on a robotic end-effector to enable learning and executing adaptive grasping based on material properties. In the acquisition platform, we position the RF sensor on a height-adjustable stand such that the antenna faces the material’s center during data acquisition, in order to maximize signal strength and ensure that measurements capture the material’s overall properties. Meanwhile, a convex lens with 15-cm-focal-length and $\sim$2 m sensing range is used to collimate the beam. With this configuration, the system achieves a beam footprint of 1–5 cm, representing the width of the beam on the material surface. In RF sensing, a smaller beam footprint produces a more focused beam with higher energy concentration, enabling more precise measurement of local material properties and yielding richer, more discriminative representations for learning-based models, thereby supporting fine-grained material classification. The lens ensures that the beam footprint is sufficiently small to distinguish materials of varying-sized objects under typical indoor conditions, remains robust to changes in sensing distance and background materials, and provides a consistent sensing region across samples, thereby mitigating potential biases arising from differences in material plate sizes. Appendix 10.1 visualizes effective beam footprint on material plates.
4 Dataset
Fine-grained material identification is challenging due to subtle differences in subcategory materials, sensitivity of RF responses to geometric perturbations, and variability in the physical information captured across frequency bands. Therefore, rich data diversity in terms of fine-grained material categories, geometric perturbation simulation, and broad frequency band coverage is crucial for developing and evaluating algorithms applicable to indoor embodied AI scenarios. Thus, we introduce RF-MatID, the first large-scale, wide-band, and geometry-diverse RF dataset for material identification, containing 142k samples evenly split across dual-domain representations, with 71k in the frequency domain and 71k in the time domain. It provides 16 fine-grained subcategories, organized into five commonly encountered superclasses in indoor scenarios. To simulate the geometric perturbation from real-world data collection, samples in RF-MatID are acquired across a distance range of 200 mm to 2000 mm (at 50 mm intervals) and an angle range of 0° to 10° (at 1° intervals). Each sample spans a wide spectral band of 4 – 43.5 GHz, uniformly represented by 2,048 bins.
4.1 Categories of materials
RF-MatID encompasses 16 fine-grained material categories organized into five superclasses. These superclasses represent the most common materials in indoor embodied AI scenarios: (i) bricks, (ii) glass, (iii) synthetic materials, (iv) woods, (v) stones. Within each superclass, we select multiple variants that exhibit subtle physical differences, enabling a rigorous evaluation of learning-based approaches on challenging fine-grained material classification. Specifically, the fine-grained material categories include: for bricks, (a) overfired clay brick, (b) lightweight perforated brick, (c) lava brick; for glass, (d) transparent acrylic glass, (e) tempered glass, (f) white opaque acrylic glass; for synthetic materials, (g) melamine-faced chipboard, (h) mineral fiber board, (i) solid polyvinyl chloride sheet; for woods, (j) cedar sleeper, (k) luan plywood, (l) red oak plywood; and for stones, (m) permeable paving Stone, (n) agglomerated stone, (o) granite, (p) concrete. Figure 2 presents pictures of the 16 fine-grained materials, annotated with their corresponding tag identifiers for reference. The detailed material sample descriptions are provided in the appendix.
Figure 2: The visual illustration of the 16 fine-grained material categories.
4.2 Data collection and preprocessing
Realistic Data Collection
Motivated by the geometric perturbations typically encountered in data acquisition, we introduce controlled variations in both distance and incidence angle, as illustrated in Figure 1(b). The Friis transmission equation(friis1946note) indicates that, for fixed transmit power $P_{t}$, antenna gains $G_{t}$ and $G_{r}$, and wavelength $\lambda$, the received signal power $P_{r}$ decreases as the effective propagation path length $(r=2d)$ increases. Furthermore, for generally rough surfaces, the backscattered signal strength diminishes as the incidence angle $\theta_{i}$ increases relative to the incident wave, due to the cosine-dependent reduction in the backscattering coefficient $\sigma_{i}$, as described by the small perturbation method (el2004multiple). In embodied AI applications, these geometric variations reflect realistic conditions where compact UWB-mmWave sensors deployed on manipulators encounter varying hand–object distances and changing incidence angles. Such variations introduce systematic changes in the RF signal representations defined in equation 1, requiring models to learn invariance and generalize across these sensing conditions.
$P_{r}=P_{t}G_{t}G_{r}({\lambda}/{4\pi r})^{2}$ (2)
$\sigma_{i}=8k^{4}_{x}\delta^{2}cos^{4}\theta_{i}\left|\alpha_{pq}\right|^{2}W(2k_{x}sin\theta_{i}),\sigma_{i}\propto cos^{4}\theta_{i}$ (3)
Concretely, for each material, samples are systematically collected across various distances and angles configurations. Data acquisition begins at a distance of 200 mm, with an angle of 0°, where 20 samples are recorded. Each sample consists of 2048 frequency-domain bins covering the 4–43.5 GHz band. The angle is then incremented from 1° to 10°, with 10 samples collected at each angle. After completing all angles at the current distance, the distance is increased in 50 mm steps up to 2000 mm, and the same angular sampling procedure is repeated at each step. The process ensures dense coverage of the distance–angle space while increasing the sampling number at normal incidence. Appendix Figure 6 & 7 shows the effect of distance and incidence angle on the frequency signal. Other forms of realistic perturbations are also discussed in the appendix 8.4.
Domain Transformation
To provide a dual-domain representation, each frequency-domain signal is paired with a time-domain signal via inverse fast Fourier transform, capturing complementary spectral and temporal features of the material response. Concretely, the 2048 $(N)$ frequency bins spanning from 4 GHz $(f_{start})$ to 43.5 GHz $(f_{end})$ with uniform spacing $\Delta f=(f_{end}-f_{start})/(N-1)\approx 19.3$ MHz are converted to a 10240 $(N_{t})$-length time-domain signal. Prior to transformation, the frequency spectrum is multiplied by a band-pass filter and zero-padded with $\lfloor f_{start}/\Delta f\rfloor$ leading zeros to ensure correct frequency alignment. The time resolution is $\Delta t=1/(N_{t}\Delta f)$, and the time axis is computed as $z=ct/2$ with $c=3\times 10^{8}$ m/s. Both the frequency-domain and time-domain samples are saved using comma-separated values data format, i.e., “.csv” files.
Frequency-Domain Representation
For frequency-domain data, a key question arises: should each frequency bin be treated as a single complex element, or should its real and imaginary parts be represented as two separate channels, i.e., a 2D $(length\times 2)$ real-valued vector? To address this, we train two models: a deep complex network (trabelsi2017deep) designed for complex-valued inputs and a Bi-LSTM (huang2015bidirectional) with input dimension two for the dual-channel representation. Our experiments show that the dual-channel representation achieves higher classification accuracy, providing empirical evidence that phase information embedded in the real and imaginary parts can be effectively leveraged by learning-based frameworks.
Post-Processing
To improve the stability and efficiency of model learning, time-domain data is standardized to have zero mean and unit variance. For the frequency-domain data in dual-channel representation, separately normalizing the real and imaginary parts would destroy the correlation (phase) information between channels. Therefore, we perform complex whitening on the combined frequency-domain data. The data is centered by subtracting its mean. A whitening transform (koivunen1999feasibility) is then applied by multiplying with the inverse square root of the covariance matrix, obtained via eigen-decomposition. The resulting signal has zero mean, unit variance, and an identity covariance matrix, ensuring that the real and imaginary components are scaled uniformly while preserving the phase information, with its magnitude normalized.
4.3 Intended Uses
RF-MatID is designed to facilitate a broad spectrum of research and applications, especially in the field of embodied AI. As the first publicly available RF material identification dataset and benchmark with diverse protocols and settings, it enables fair comparisons and accelerates the advancement of material identification algorithms. By providing a standardized dataset aligned with calibrated radar systems, RF-MatID enables the evaluation of models’ transferability to real-world application scenarios. For instance, models can be trained on the 24 GHz sub-band in RF-MatID and then evaluated on radar signals captured by embodied agents operating in practical indoor scenarios, enabling assessment of their zero-shot or few-shot transferability. RF-MatID further facilitates research on domain adaptation and generalization through cross-domain evaluation settings, advancing algorithm robustness under diverse domains and environmental conditions. From a modality perspective, RF-MatID introduces a compact, low-cost sensing platform operating in the UWB–mmWave spectrum. The platform is specifically customized for indoor embodied AI applications, as it can be integrated onto a robot’s end-effector to enable fine-grained local material characterization. This capability supports material aware manipulation and affordance driven workflows, such as selecting grasp strategies according to material compliance, adjusting contact forces based on surface hardness, and enabling downstream policies that rely on material grounded affordance cues, for example graspable, cuttable, or pourable objects. Moreover, the sensing platform can serve as a complementary modality within multimodal learning frameworks, providing materials’ electromagnetic characteristics that enrich the information available for embodied perception. For example, when combined with vision-based perception pipelines in indoor scene understanding, the UWB-mmWave signals provide fine grained local material characteristics that can be further mapped to material grounded affordances of individual objects. When fused with visual cues that capture spatial structure and object geometry in cluttered environments, this enables embodied agents to achieve more comprehensive scene understanding across multiple dimensions, ultimately supporting more reliable reasoning and action in complex indoor settings.
5 Benchmark and Evaluation
In this section, we present the key benchmark configurations, the evaluation metrics, the selected deep learning models, and the baseline model design for material identification based on our proposed RF-MatID Dataset. We further evaluate experimental results to highlight the limitations of each benchmark model and demonstrate the applicability of learning-based approaches for RF-based material identification across diverse real-world scenarios.
5.1 Benchmark Setup
Frequency Protocol
We define five frequency-band protocols for RF-based material identification, capturing both physical distinctions (centimeter vs. millimeter waves) and region-specific regulatory constraints. Protocol 1 (P1) spans the full spectrum from 4–43.5 GHz. Protocol 2 (P2) focuses on millimeter-wave analysis, covering 30–43.5 GHz, while Protocol 3 (P3) targets centimeter-wave analysis, spanning 4–30 GHz. Pioneeringly, we also consider the practical feasibility of RF-based material identification under legal frequency regulations in major global economies. Protocol 4 (P4) covers frequency bands permitted for commercial RF sensor development in the United States, and Protocol 5 (P5) covers legally allowed bands in China. The details are listed in the appendix 4.
MLP ResNet-50 BiLSTM Trans-former TimesNet LSTM-ResNet ConvNeXt DINOv3 Material-ID AirTac Baseline
Protocol 1 (4.0-43.5 GHz)
S1 - 99.19 98.85 86.33 91.76 99.66 99.84 99.51 99.28 96.81 99.77 99.57
S2 mod1 84.47 97.17 80.61 84.66 82.87 97.12 81.70 79.10 95.67 91.36 86.62
mod2 73.49 48.90 50.76 40.49 68.58 49.95 65.74 64.19 71.59 86.95 69.47
mod3 76.46 83.16 69.85 79.16 59.93 71.00 66.07 63.52 72.37 65.41 74.09
S3 mod1 99.08 99.16 86.78 87.13 98.59 99.69 99.45 98.85 97.63 98.12 98.89
mod2 90.59 59.48 77.05 53.38 78.99 78.13 85.90 86.27 51.80 76.60 85.23
mod3 96.39 90.42 86.50 82.87 75.50 80.37 94.84 95.21 70.54 75.23 94.21
Protocol 2 (30.0-43.5 GHz)
S1 - 98.11 99.82 87.26 92.61 94.78 99.80 99.34 98.56 93.53 91.13 99.47
S2 mod1 88.35 95.86 83.75 89.35 84.14 96.82 87.29 85.13 93.20 91.16 86.87
mod2 69.19 58.82 53.31 56.80 57.96 50.76 68.90 62.65 72.43 71.91 62.31
mod3 79.27 82.12 77.52 79.64 75.15 77.02 81.28 66.81 75.36 66.89 79.81
S3 mod1 96.64 95.88 83.96 88.61 93.66 99.36 98.03 95.71 86.86 88.51 97.69
mod2 84.41 51.31 61.45 57.44 67.17 65.15 75.90 73.31 53.09 69.61 75.17
mod3 93.75 94.78 85.70 88.34 87.20 97.59 95.44 94.15 78.49 81.45 96.30
Protocol 3 (4.0-30.0 GHz)
S1 - 99.52 99.82 89.87 88.22 99.64 99.78 99.35 98.53 97.32 98.76 99.47
S2 mod1 81.57 95.01 78.59 84.22 78.22 95.92 79.45 67.65 97.55 88.84 78.18
mod2 62.81 38.45 52.25 42.16 64.58 55.18 65.41 65.28 62.83 73.84 60.07
mod3 65.93 65.29 62.14 76.97 59.50 64.34 53.19 61.68 58.59 54.64 62.63
S3 mod1 99.33 99.31 89.48 90.14 98.56 99.51 99.30 98.30 98.09 95.51 99.22
mod2 89.79 79.20 76.28 62.04 86.26 78.46 90.28 90.63 82.40 87.11 87.17
mod3 95.32 81.63 87.29 74.63 72.29 79.34 93.50 93.23 65.75 62.77 93.47
Number of Model Parameters (M)
67.20 15.98 0.53 0.20 1.34 4.20 28.06 28.05 0.94 1.14 16.27
Table 2: Comprehensive benchmark of deep learning model end-to-end material classification performance on RF-MatID dataset. Accuracy is evaluated and shown in percentage (%) values. Bold indicates the best performance, while underlined values denote the second- and third-best results.
Data Splits
To evaluate the robustness of learning-based approaches under diverse real-world conditions, we define three primary data split settings with seven sub-modes. Setting 1 (S1: Random Split) randomly partitions all RF signal samples into training and testing sets at a 7:3 ratio. Setting 2 (S2: Cross-Distance Split) partitions the dataset by distance to simulate distance-domain distribution shifts. Three sub-modes are defined: (i) S2-1, where 11 uniformly spaced distances (out of 37) are used as the test set; (ii) S2-2, where the 11 closest distances are used for testing; and (iii) S2-3, where the 11 farthest distances are used for testing. Setting 3 (S3: Cross-Angle Split) partitions the dataset by incidence angle to simulate angle-domain distribution shifts, also with three sub-modes: (i) S3-1, where 3 uniformly spaced angles (out of 11) are used as the test set; (ii) S3-2, where the 3 smallest angles are used for testing; and (iii) S3-3, where the 3 largest angles are used for testing. In all of the benchmark results tables, we use the term “mod” as an abbreviation for data split modes.
Category Division
To meet various application requirements, we evaluate model performance under three category divisions. The fine-grained division treats all 16 materials as independent classes, regardless of their higher-level grouping. The superclass division groups all subclasses under each superclass into a single category (e.g., cedar sleeper, luan plywood, and red oak plywood are unified as “wood”). The subclass division assesses model accuracy in constrained, fine-grained tasks; specifically, we focus on the four subclasses under the “stone” category for analysis.
Evaluation Metrics
In our benchmark, Accuracy is reported in all experimental analyses. To address the limitation that accuracy may be dominated by majority classes, we also report the Macro F1-score alongside accuracy when presenting baseline results under different task configurations. We further include precision and recall, with the detailed results reported in the appendix 7.
Protocols Protocol 1 Protocol 2 Protocol 3 Protocol 4 Protocol 5
Settings Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
Fine- Grained Division S1 - $99.57$ $99.57$ $99.47$ $99.47$ $99.47$ $99.47$ $99.53$ $99.53$ $99.41$ $99.41$
S2 mod1 $86.62$ $86.66$ $86.87$ $86.82$ $78.18$ $78.21$ $87.35$ $87.26$ $84.82$ $84.65$
mod2 $69.47$ $68.46$ $62.31$ $62.12$ $60.07$ $59.81$ $66.27$ $65.89$ $67.18$ $66.45$
mod3 $74.09$ $73.07$ $79.81$ $79.22$ $62.63$ $62.67$ $78.40$ $77.46$ $72.04$ $70.63$
S3 mod1 $98.89$ $98.89$ $97.69$ $97.69$ $99.22$ $99.22$ $98.42$ $98.42$ $99.37$ $99.38$
mod2 $85.23$ $85.12$ $75.17$ $75.10$ $87.17$ $86.96$ $74.94$ $74.51$ $85.11$ $84.87$
mod3 $94.21$ $94.16$ $96.30$ $96.27$ $93.47$ $93.44$ $96.65$ $96.63$ $95.95$ $95.90$
Superclass Division S1 - $99.82$ $99.82$ $99.78$ $99.78$ $99.82$ $99.82$ $99.85$ $99.85$ $99.87$ $99.87$
S2 mod1 $91.15$ $90.85$ $94.58$ $94.43$ $91.15$ $90.85$ $93.50$ $93.30$ $94.00$ $93.98$
mod2 $87.24$ $86.58$ $88.54$ $87.98$ $86.05$ $85.86$ $84.69$ $84.09$ $85.19$ $84.61$
mod3 $87.23$ $87.50$ $89.54$ $89.51$ $86.48$ $86.46$ $89.31$ $89.67$ $87.34$ $87.61$
S3 mod1 $99.71$ $99.69$ $99.34$ $99.30$ $99.44$ $99.40$ $99.18$ $99.15$ $99.69$ $99.68$
mod2 $91.26$ $90.73$ $89.08$ $88.64$ $94.29$ $93.96$ $85.80$ $84.74$ $91.68$ $91.09$
mod3 $98.11$ $98.10$ $99.06$ $99.02$ $97.94$ $97.87$ $98.99$ $98.96$ $98.02$ $98.01$
Subclass Division S1 - $99.61$ $99.61$ $99.72$ $99.72$ $99.68$ $99.68$ $99.66$ $99.66$ $99.72$ $99.72$
S2 mod1 $98.77$ $98.77$ $96.36$ $96.36$ $95.55$ $95.51$ $96.48$ $96.45$ $98.05$ $98.05$
mod2 $85.76$ $85.54$ $80.57$ $80.43$ $82.16$ $81.90$ $86.70$ $86.54$ $90.40$ $90.35$
mod3 $85.81$ $85.41$ $85.80$ $85.54$ $81.67$ $81.63$ $78.75$ $76.40$ $84.15$ $83.59$
S3 mod1 $99.17$ $99.17$ $98.81$ $98.80$ $99.10$ $99.10$ $99.55$ $99.55$ $99.05$ $99.05$
mod2 $95.63$ $95.62$ $88.04$ $87.97$ $95.37$ $95.37$ $87.57$ $87.59$ $95.39$ $95.36$
mod3 $98.29$ $98.29$ $97.75$ $97.74$ $98.87$ $98.87$ $99.08$ $99.08$ $98.18$ $98.17$
Table 3: Baseline model performance under various category divisions, data split settings, and protocols. Accuracy and macro F1 score are evaluated and shown in percentage (%) values.
Benchmark Models and Baseline Design
Considering the sequential dependencies across frequency bins as well as the spatial features along the frequency and channel dimensions, we benchmark a diverse set of models commonly used in computer vision, natural language processing, time-series, and RF-sensing research. These include Multilayer Perceptron (MLP) (gardner1998artificial), ResNet-50 (he2016deep), Bidirectional LSTM (BiLSTM) (huang2015bidirectional), Vanilla Transformer (vaswani2017attention), TimesNet (wu2022timesnet), Material-ID (chen2025material), AirTac (zhang2024airtac), a hybrid LSTM–ResNet model (choi2018short), ConvNeXt (liu2022convnet), and the recent DINOv3 (simeoni2025dinov3). We also introduce a simple yet robust baseline model that leverages frequency-aware positional encoding to preserve global consistency. Parallel extractors independently capture spatial and temporal features, which are then integrated into class probabilities via an MLP fusion module. The baseline model achieves an average accuracy of 85% in all experimental configurations, while the other models perform at approximately 80%. Detailed model implementations are provided in the appendix 9.4.
5.2 Results and Analytics
Domain Comparison
Table 4 summarizes material identification results on time- and frequency-domain signals. Under Protocol 1 for fine-grained classification, LSTM-ResNet achieves comparable performance on time-domain data to the baseline model on frequency-domain data. Given the additional effort to convert frequency signals into the time domain, and significantly higher computational complexity when using 10,240-length time-domain data, we conclude that directly learning from raw frequency-domain data is more optimized and efficient.
Benchmark Across Models
To benchmark model performance and highlight their strengths and weaknesses in material identification, we evaluate all split settings under P1–3 using the fine-grained division, as shown in Table 2. The MLP shows consistently strong performance across configurations, but it has the largest parameter size and requires careful redesign of intermediate embedding dimensions to adapt to varying protocols. ResNet-50 and the LSTM-ResNet excel under mild domain shifts (S1, S2-1, S3-1) but degrade sharply under severe shifts (S2-2/3, S3-3). SOTA vision-based models such as DINOv3 and ConvNeXt perform competitively in challenging scenarios like S3-3. However, the low resolution of the RF data hampers the stability of model convergence, leading to suboptimal results in simpler settings. Sequence-oriented models (Vanilla Transformer, BiLSTM, TimesNet) underperform in most settings and exhibit particular vulnerability to domain shifts. RF-sensing models generally perform well, but still exhibit notable performance drops under certain out-of-distribution conditions. The baseline model demonstrates robustness and competitive performance in most cases.
Time-Domain Freq-Domain
Acc F1 Acc F1
S1 - 96.95 96.95 99.57 99.57
S2 mod1 85.64 85.29 86.62 86.66
mod2 71.61 69.60 69.47 68.46
mod3 74.97 74.32 74.09 73.07
S3 mod1 99.65 99.65 98.89 98.89
mod2 94.43 94.42 85.23 85.12
mod3 93.70 93.60 94.21 94.16
Table 4: Comparison between material identification performance on time- and frequency-domain signals.
Figure 3: Preliminary experiments on consecutive sub-bands with various bandwidths. Accuracy in percentage value is applied for performance evaluation.
Quantitative Baseline Results
By evaluating the baseline model across all three category divisions, seven split settings, and five protocols, we draw several insights from Table 3. RF-based approaches achieve an average material identification accuracy of 96.83% under S1 across all protocols and divisions. However, performance is significantly impacted by domain shifts, with distance shifts causing an average 24.17% drop and angle shifts a 11.41% drop. We further observe that millimeter-wave signals (P2) are more robust to distance variations, whereas centimeter-wave signals (P3) better tolerate changes in incidence angle. Interestingly, country-specific legal bands (P4 and P5) achieve performance comparable to the full spectrum (P1), demonstrating the feasibility of RF-based material identification under legal constraints without significant accuracy degradation.
Experiments on Consecutive Sub-bands with Various Bandwidths
Band analysis offers valuable guidance for selecting optimal operating frequencies and designing tailored algorithms for specific applications. In this work, we present a preliminary exploration: as shown in Figure 3, RF signals across different bandwidths generally achieve high accuracy ($95\%$) on S1. However, noticeable drops appear at certain ranges (e.g., around 10 GHz and in the higher band 35–43.5 GHz), offering useful insights for selecting effective frequency ranges in RF-based material identification.
Discussions
By analyzing the benchmark results, we highlight several challenges of learning-based material identification approaches and discuss possible solutions. Training with a standard classification loss captures only data-driven correlations, neglecting physically consistent features, which leads to instability under domain shifts and unconstrained intermediate features. Incorporating physical constraints (e.g., via PINNs) guides meaningful feature learning, improving interpretability and robustness. Additionally, standard training often overfits the source domain, limiting generalization. Domain adaptation and generalization techniques address this by aligning features or adapting parameters, enhancing cross-domain transfer and overall robustness.
6 Limitations and Conclusion
For future improvements, we identify the following limitations in the RF-MatID dataset: First, the dataset could be expanded to include richer material variability; it does not yet cover complexities such as multi-layer composite materials or significant thickness-induced signal shifts. Second, the dataset lacks diverse environmental contexts. Future work should incorporate more complex settings, such as cluttered backgrounds or large open spaces, to account for additional multipath effects and occlusion-induced biases found in real-world scenarios. Third, the RF data is sparsely sampled across the broad frequency spectrum, and the raw complex-valued signals have low feature dimensionality. To address these issues in the next-generation dataset, we will introduce variables such as material thickness and area to enrich the diversity of material samples, expand to real-world data collection in outdoor production and construction scenarios, increase the sampling rate in application-relevant frequency bands, and leverage classical signal processing techniques to expand and enrich signal feature dimensions.
In this paper, we present the first open-source, large-scale, wide-band, and geometry-diverse RF dataset for fine-grained material identification, covering 16 fine-grained categories across the 4–43.5 GHz band with controlled variations in angle and distance, and providing both time- and frequency-domain representations. We further show that raw frequency-domain signals can be effectively leveraged by deep learning models without additional domain transformations, and we evaluate their applicability under versatile protocols. Finally, by benchmarking state-of-the-art models and systematically assessing their robustness to out-of-distribution shifts, our work provides a critical foundation for developing more reliable and physically grounded RF sensing systems.

### v2

@@
 1 Introduction
-Material identification is the task of identifying an object’s physical category (e.g., metal, plastic) from its intrinsic properties, which is fundamental for autonomous systems, particularly in embodied intelligence.
+Material identification is the task of identifying an object’s physical category (e.g., metal, plastic) from its intrinsic properties, which is fundamental for physical artificial intelligence (Physical AI), particularly in embodied intelligence.
 Embodied intelligence arises from an agent’s physical interaction with the environment, enabling it to perceive, reason, and adapt.
@@
 There are no publicly accessible large-scale datasets, hindering reproducibility and fair benchmarking across algorithms.
-In addition, commercial off-the-shelf (COTS) sensors offer limited and fragmented frequency coverage, hindering systematic evaluation of material identification across the RF spectrum, which is crucial for selecting optimal operational bands in various applications.
+In addition, commercial off-the-shelf (COTS) sensors offer limited and fragmented frequency coverage (yang2024mm), hindering systematic evaluation of material identification across the RF spectrum, which is crucial for selecting optimal operational bands in various applications.
 Finally, most studies are conducted in controlled laboratory settings and rarely incorporate perturbations such as variations in sensor–object geometry (e.g., incidence angle, stand-off-distance), leaving open questions regarding robustness and scalability in deployment scenarios.
-Thus, we present RF-MatID, a large-scale, wide-band, and geometry-diverse RF dataset designed to advance fine-grained material identification and enable the development of more robust algorithms.
+Thus, we present RF-MatID, a large-scale, wide-band, and geometry-diverse RF dataset designed to advance fine-grained material identification and enable the development of more robust Physical AI algorithms.
 RF-MatID encompasses 16 carefully curated fine-grained categories, systematically derived from 5 superclasses to capture subtle intra-class variations.
@@
 The modality enables applications such as through-wall sensing via strong low-frequency penetration, but the reliance on physical tags greatly complicates data collection.
-Wi-Fi-based datasets primarily leverage the 2.4 GHz and 5 GHz frequency bands, and the target object is typically placed between a Wi-Fi transmitter and receiver (feng2019wimi; shi2021environment).
+Wi-Fi-based datasets primarily leverage the 2.4 GHz and 5 GHz frequency bands, and the target object is typically placed between a Wi-Fi transmitter and receiver (yang2023sensefi; feng2019wimi; shi2021environment).
 The ubiquity of Wi-Fi devices facilitates convenient data collection, but the narrow bandwidth and distortions induced by hardware imperfections constrain the reliability of signals.
@@
 The process ensures dense coverage of the distance–angle space while increasing the sampling number at normal incidence.
-Appendix Figure 6 & 7 shows the effect of distance and incidence angle on the frequency signal.
+Appendix Figure 6 & 7 show the effect of distance and incidence angle on the frequency signal.
 Other forms of realistic perturbations are also discussed in the appendix 8.4.
@@
 This capability supports material aware manipulation and affordance driven workflows, such as selecting grasp strategies according to material compliance, adjusting contact forces based on surface hardness, and enabling downstream policies that rely on material grounded affordance cues, for example graspable, cuttable, or pourable objects.
-Moreover, the sensing platform can serve as a complementary modality within multimodal learning frameworks, providing materials’ electromagnetic characteristics that enrich the information available for embodied perception.
+Moreover, the sensing platform can serve as a complementary modality within multimodal learning frameworks, providing materials’ electromagnetic characteristics that enrich the information available for embodied perception (chen2025xfi; AN2026101429).
 For example, when combined with vision-based perception pipelines in indoor scene understanding, the UWB-mmWave signals provide fine grained local material characteristics that can be further mapped to material grounded affordances of individual objects.
-When fused with visual cues that capture spatial structure and object geometry in cluttered environments, this enables embodied agents to achieve more comprehensive scene understanding across multiple dimensions, ultimately supporting more reliable reasoning and action in complex indoor settings.
+When fused with visual cues that capture spatial structure and object geometry in cluttered environments, this enables embodied agents to achieve comprehensive scene understanding across multiple dimensions, ultimately supporting more reliable reasoning and action in complex indoor settings.
 5 Benchmark and Evaluation
@@
 To address the limitation that accuracy may be dominated by majority classes, we also report the Macro F1-score alongside accuracy when presenting baseline results under different task configurations.
-We further include precision and recall, with the detailed results reported in the appendix 7.
+We further include precision and recall, with the detailed results reported in the appendix Table 7.
 Protocols Protocol 1 Protocol 2 Protocol 3 Protocol 4 Protocol 5